17) Given the quadratic function g(x)= -x2 - 4x + 4 a) Express the quadratic function in standard forn b) Find its vertex.
Answer #1

comma g00) = -4?_4% +4 =-[x?+ 2x2x24 +22-22]+4 --I OC+2)2 4] + 4 =-02 +232 +4+4 9100) = -(x+2)+8 which is equivalent of Stand